# Forecasting Local Ionospheric Parameters Using Transformers

### v1

@@ 4.2 Model Forecast Impacts
Numerical ray tracing offers additional insight into the performance of our ionospheric forecast model over simply driving down the MSE as compared to climatology. One of the primary use cases for ionospheric forecasting is to simulate radio propagation for long-range radar and communications systems. The ray tracing model we use is based on the Jones Stephenson 3D magnetoionic model that is prevalent throughout the space weather community [57]. We numerically solve for the ray paths using EDPs constructed from our predicted parameters, specifically, foF2 and hmF2. The ray paths allow us to determine the distances a radio wave at each frequency will propagate down range. In doing so, we may illustrate how an increase in the prediction accuracy of these parameters and their uncertainty quantification translates directly into measurable improvements for operational users.
@@
Figure 21 shows frequency-range plots for a hypothetical radio transmitter for which the sounding station GU513 (Guam, USA) is roughly at the midpoint of its propagation path. We randomly selected a time from our test period and reconstructed EDPs using parameters from the sounder measurements, our LIFT model, and the IRI climatology and then traced rays at frequencies from 3-30MHz through each EDP to give full coverage of the HF radio band. We see in Figure 21 that the forecasted frequency-range plot more closely matches that of the true profile. Similar plots are regularly used to help inform the operational space weather community as to what frequencies should be used to propagate out certain distances, wherein the leading edge of these heatmaps shows the maximum usable frequency for a given range. Regarding the leading edge curve generated by the LIFT model’s median prediction, it represents a significant improvement over IRI, even though IRI is being parameterized with the F10.7cm solar flux index from the timestamp of the last input to the encoder module.
@@
Figure 21: Simulated frequency-range plots using vertical electron density profiles generated from the observed foF2 and hmF2 (True), the LIFT model (Forecasted), and IRI predicted (IRI). The date for this simulation was randomly selected but constrained to a period in 2019 of low solar activity.
@@
However, the leading edge curve from the median prediction is only the point forecast. Using the LIFT model CIs for foF2 and hmF2, we may generate additional leading edge curves and produce uncertainty quantification in these downstream propagation tools as well. While Equation 12 represents the $90\%$-CI for both foF2 and hmF2 simultaneously, we may also produce different permutations of the quantile outputs from LIFT for each parameter. Figure 22 illustrates the leading edge curves for these permutations of the estimated LIFT quantiles. These curves reveal the inverse relationship between foF2 and hmF2, where the true leading edge will almost always be bracketed on the left by the combination ($5^{th}$-percentile foF2, $95^{th}$-percentile hmF2) and on the right by ($95^{th}$-percentile foF2, $5^{th}$-percentile hmF2). These represent the extremes in our forecast, and while their purpose may remain illustrative in this work, they further motivate the need for multivariate uncertainty quantification in future ionospheric forecasting models.
@@
Figure 22: Simulated leading edges of frequency-range plots using vertical electron density profiles generated from the observed foF2 and hmF2 (True), IRI predicted (IRI), and the upper and lower 5th percentiles from the LIFT model (Forecasted). The date for this simulation was randomly selected but constrained to a period in 2019 of low solar activity.
Figures 21 and 22 only illustrate a single time stamp from the 24-hour LIFT model forecast. By generating the leading edge curves in Figure 22 for every time step in the forecast, we show which combination of quantiles most accurately matched the leading edge derived from the sounder profile as a function of time and range. Figure 23 shows this as a heatmap, and Figure 24 provides the histogram for how many range-time bins each combination was the best match of the true leading edge. The histogram indicates that this forecast period was particularly well matched by the median LIFT output ($50^{th},50^{th}$). See Appendices A and B for additional frequency-range and leading edge plots for periods of low and high solar activity, respectively.
@@
Figure 24: Histogram reporting which quantile combination of foF2 and hmF2 parameters from the LIFT forecast produced the most accurate HF propagation in Guam, USA during a periods of low solar activity.
Although the purpose of these plots in this work is purely qualitative, future efforts to improve LIFT models may seek to incorporate these types of propagation results into the calibration of the model uncertainty. While numerical ray tracing is a computationally expensive process to incorporate into the training process, it is also possible that substituting more efficient approximations could also be useful in performing model calibration.
@@ 5 Summary and Conclusions
We have demonstrated that transformer models can be used in a pseudo-data assimilation fashion that fuses multiple space weather data sources with naive predictors. In this method, we treat both the simple linear autoregression and the IRI climatology as naive forecasts. While the transformer makes including exogenous covariates very straightforward, it is also a natural choice for quantile forecasting and uncertainty quantification due to its ability to translate long-range influences in the data into confidence regions that are tailored to each data segment rather than relying on assumptions about the underlying distributions of the errors. The other advantage of transformers for space weather forecasting is their efficiency in processing long time series with relatively high cadence in comparison to conventional recurrent neural networks (RNNs) such as long short-term memory (LSTM) networks. We also establish how the predicted confidence regions for these parameters translate directly into real-world applications for more robust radio propagation assessments and planning through ray traced frequency-range plots.
@@
Although the LIFT model approach addresses several challenges in modern space weather forecasting, it is nevertheless limited to local forecasts and may not provide the most robust predictions under severely disturbed conditions. Future iterations of this approach will therefore focus on extensions to global grids and better calibration of uncertainty, especially in the tails of the distributions and during perturbed conditions. It is important to reiterate that the data used in this study were minimally processed from the original GIRO dataset, only using the ARTIST5 confidence scores to remove poorly autoscaled data. This was intentional and meant to demonstrate that we could train a reasonably well-calibrated system using data as one would expect it to find it in real-time from the GIRO repository. This is in contrast to many other data-driven models built from only exquisite, manually scaled EDPs. Finally, transformers require large amounts of data to train on, and time series prediction in particular is much easier when data are contiguous. This is not the case for the other common parameters for the ionospheric profile that specify the F1- and E- regions, but future work is required to include these types of data in the output of this model as they only intermittently available, i.e. only measurable during certain hours of the day or during certain space weather conditions.
6 Availability Statement
@@
The model code is openly available at https://github.com/JayLago/LIFT. The ionosonde data may be found at https://giro.uml.edu/didbase/. The SSN, Kp, and F10.7 indices were obtain from the GFZ German Research Centre for Geosciences, ftp://datapub.gfz-potsdam.de/download/10.5880.Kp.0001. Storm index, Dst, was obtained through the WDC for Geomagnetism, Kyoto http://wdc.kugi.kyoto-u.ac.jp/wdc/Sec3.html.
@@ Acknowledgments
The authors would like to thank Dr. Victoriya Forsythe for her assistance and input with running the PyIRI model, the code for which can be found at [52].The authors would like to also thank Dr. Ivan Galkin from the LGDC for the data they continue to collect and make available. We would also like to acknowledge the authors of [41, 26] for the code they made available and which initially formed a basis for constructing our model.
Appendix A Frequency Versus Range Plots from the LIFT Model During Low Solar Activity
The following presents visualizations of the performance of the LIFT model when used with numerical ray tracing to estimate the frequency versus range curves within the HF radio band during low solar activity. These results were generated for a randomly selected sounding station and time from the test set. Presented are the predicted frequency-range heatmaps and leading edge curves at $+0,+3,+6,+9,+12,+15,+18$ and $+21$ hours. The station in Learmonth, Australia, was selected from the test set and a date was randomly chosen from the period 2019-7-1 00:00Z to 2019-12-31 00:00Z. This occurred roughly during the last solar minimum. Figures 25 to 32 illustrate the frequency-range plots and affiliated leading edges. Figure 33 summarizes which forecasted uncertainty ranges for different levels of CI for each parameter (foF2, hmF2) produced the most accurate estimates of the leading edge as a function of time and range.
Figure 25: Forecast +0 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 26: Forecast +3 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 27: Forecast +6 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 28: Forecast +9 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 29: Forecast +12 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 30: Forecast +15 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 31: Forecast +18 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 32: Forecast +21 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 33: Best forecasted quantile pairings (foF2, hmF2) at the given timestamp for station LM42B in Learmonth, Australia located at -21.80 latitude, 114.10 longitude.
Appendix B Frequency Versus Range Plots from the LIFT Model During High Solar Activity
The following presents visualizations of the performance of the LIFT model when used with numerical ray tracing to estimate the frequency versus range curves within the HF radio band during high solar activity. These results were generated for a randomly selected sounding station and time from the test set. Presented are the predicted frequency-range heatmaps and leading edge curves at $+0,+3,+6,+9,+12,+15,+18$ and $+21$ hours. The station in Learmonth, Australia, was selected from the test set and a date was randomly chosen from the period 2013-12-1 00:00Z to 2014-5-30 00:00Z. This occurred roughly during the last solar maximum. Figures 25 to 32 illustrate the frequency-range plots and affiliated leading edges. Figure 33 summarizes which forecasted uncertainty ranges for different levels of CI for each parameter (foF2, hmF2) produced the most accurate estimates of the leading edge as a function of time and range.
Figure 34: Forecast +0 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 35: Forecast +3 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 36: Forecast +6 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 37: Forecast +9 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 38: Forecast +12 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 39: Forecast +15 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 40: Forecast +18 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 41: Forecast +21 hours. (Top) Frequency vs. range plot using EDPs derived from foF2 and hmF2 parametres from: the sounder measurement (Ionosonde), LIFT model median, and IRI. (Bottom) The leading edge of the frequency vs. range plots using: the sounder measurement (sounder), IRI climatology, and all combinations of the LIFT predicted quantiles.
Figure 42: Best forecasted quantile pairings (foF2, hmF2) at the given timestamp for station LM42B in Learmonth, Australia located at -21.80 latitude, 114.10 longitude.

### v2

@@ 4.2 Model Forecast Impacts
A common application for improved local ionospheric parameter forecasts is in propagation analysis for long-range radio communications systems. Propagation models are only as good as the environmental specification provided to them, and so in this section we illustrate how the LIFT model can offer improved frequency vs range prediction for high-frequency (HF) radio propagation over PyIRI that includes uncertainty quantification. To do this, we use a numerical ray tracing model that is based on the original Jones Stephenson 3D magnetoionic HF propagation code (?, ?). We numerically solve for the ray paths using electron density profiles (EDPs) constructed from LIFT’s predicted parameters, specifically, foF2 and hmF2. The ray paths then allow us to determine the distances a radio wave at each frequency will propagate down range. In doing so, we may illustrate how an increase in the prediction accuracy of these parameters and their uncertainty quantification translates directly into measurable improvements for operational users.
@@
Figure 22 shows frequency-range plots for a hypothetical radio transmitter for which the sounding station GU513 (Guam, USA) is roughly at the midpoint of its propagation path. We randomly selected a time from our test period and reconstructed EDPs using parameters from the sounder measurements, our LIFT model, and PyIRI, and then traced rays at frequencies from 3-30MHz through each EDP to give full coverage of the HF radio band. We see in Figure 22 that the forecasted frequency-range plot more closely matches that of the true profile. Similar plots are regularly used to help inform the operational space weather community as to what frequencies should be used to propagate out certain distances, wherein the leading edge of these heatmaps shows the maximum usable frequency for a given range. Regarding the leading edge curve generated by the LIFT model’s median prediction, it represents a significant improvement over PyIRI, even though PyIRI is being parameterized with the F10.7cm solar flux index from the timestamp of the last input to the encoder module.
@@
Figure 22: Simulated frequency-range plots using vertical electron density profiles generated from the observed foF2 and hmF2 (True), the LIFT model (Forecasted), and PyIRI predicted (IRI). The date for this simulation was randomly selected but constrained to a period in 2019 of low solar activity.
@@
However, the leading edge curve from the median prediction is only the point forecast. Using the LIFT model CIs for foF2 and hmF2, we may generate additional leading edge curves and produce uncertainty quantification in these downstream propagation tools as well. While Equation 12 represents the $90\%$-CI for both foF2 and hmF2 simultaneously, we may also produce different permutations of the quantile outputs from LIFT for each parameter. Figure 23 illustrates the leading edge curves for these permutations of the estimated LIFT quantiles. These curves reveal the inverse relationship between foF2 and hmF2, where the true leading edge will almost always be bracketed on the left by the combination ($5^{th}$-percentile foF2, $95^{th}$-percentile hmF2) and on the right by ($95^{th}$-percentile foF2, $5^{th}$-percentile hmF2). These represent the extremes in our forecast, and while their purpose remains purely illustrative in this work, they further motivate the need for multivariate uncertainty quantification in future ionospheric forecasting models used for propagation analysis.
@@
Figure 23: Simulated leading edges of frequency-range plots using vertical electron density profiles generated from the observed foF2 and hmF2 (True), PyIRI predicted (IRI), and the upper and lower 5th percentiles from the LIFT model (Forecasted). The date for this simulation was randomly selected but constrained to a period in 2019 of low solar activity.
Figures 22 and 23 only illustrate a single time step from the 24-hour LIFT model forecast. By generating the leading edge curves in Figure 23 for every time step in the forecast, we show which combination of quantiles most accurately matched the leading edge derived from the sounder profile as a function of time and range. Figure 24 shows these quantile pairings as a heatmap over the entire 24-hour forecast, and Figure 25 provides the histogram for how many range-time bins each combination was the best match of the true leading edge. The histogram indicates that this forecast period was particularly well matched by the median LIFT output ($50^{th},50^{th}$). However, Figures 26 and 27 demonstrate that the median prediction is not always the best. This particular period in 2014 captures a moderate geomagnetic storm (Dst $\leq$ -50 nT). As a result, the LIFT model ($5^{th},50^{th}$) quantile prediction more closely matched the true leading edge for a station in Learmonth, Australia. See Supporting Information for additional frequency-range and leading edge plots during this period. This indicates that, while the median was a poor predictor during the storm, the lower bound of the foF2 forecast was still able to generate a leading edge that covered the observations. While this helps illustrate the importance of uncertainty quantification in ionospheric forecasts, further development of these types of tools will need to be the focus of future work. Improvements to the LIFT model may even seek to incorporate propagation results into the calibration of the model uncertainty. Numerical ray tracing is computationally very expensive, however, it is possible that substituting more efficient approximations such as surrogate models could prove useful in performing model calibration.
@@
Figure 25: Histogram reporting which quantile combination of foF2 and hmF2 parameters from the LIFT forecast produced the most accurate HF propagation in Guam, USA during a period of low solar activity.
Figure 26: Performance of LIFT quantile forecast in Learmonth, Australia during a period of high solar activity and during a moderate geomagnetic storm (Dst $\leq$ -50 nT).
Figure 27: Histogram reporting which quantile combination of foF2 and hmF2 parameters from the LIFT forecast produced the most accurate HF propagation in Learmonth, Australia during a period of high solar activity and during a moderate geomagnetic storm (Dst $\leq$ -50 nT).
@@ 5 Summary and Conclusions
We have presented LIFT, a Local Ionospheric Forecast Transformer that jointly predicts foF2, hmF2, and TEC with calibrated uncertainty quantification. To our knowledge, LIFT is the first transformer-based ionospheric model to combine joint multi-parameter local forecasting, multi-quantile uncertainty bounds, a hybrid linear-transformer architecture, and rigorous spatial generalization testing on geographically held-out stations. The model operates in a pseudo-data assimilation fashion that fuses multiple space weather data sources with naïve predictors, treating both the simple linear autoregression and the PyIRI climatology as naïve forecasts.
@@
While the transformer makes including exogenous covariates very straightforward, it is also a natural choice for quantile forecasting and uncertainty quantification due to its ability to translate long-range influences in the data into confidence regions that are tailored to each data segment rather than relying on assumptions about the underlying distributions of the errors. The other advantage of transformers for space weather forecasting is their efficiency in processing long time series with relatively high cadence in comparison to conventional recurrent neural networks (RNNs) such as long short-term memory (LSTM) networks. We also establish how the predicted confidence regions for these parameters translate directly into real-world applications for more robust radio propagation assessments and planning through ray traced frequency-range plots.
@@
Although the LIFT model approach addresses several challenges in modern space weather forecasting, it is nevertheless limited to local forecasts and may not provide the most robust predictions under severely disturbed conditions. Future iterations of this approach will therefore focus on extensions to global grids and better calibration of uncertainty, especially in the tails of the distributions and during perturbed conditions. It is important to reiterate that the data used in this study were minimally processed from the original GIRO dataset, only using the ARTIST5 confidence scores to remove poorly autoscaled data. This was intentional and meant to demonstrate that we could train a reasonably well-calibrated system using data as one would expect to find it in real-time from the GIRO repository. This is in contrast to many other data-driven models built from only exquisite, manually scaled EDPs. Finally, transformers require large amounts of data to train on, and time series prediction in particular is much easier when data are contiguous. This is not the case for the other common parameters for the ionospheric profile that specify the F1- and E-regions, but future work is required to include these types of data in the output of this model as they are only intermittently available, i.e. only measurable during certain hours of the day or during certain space weather conditions.
Data Availability
@@
The model code is openly available at https://github.com/JayLago/LIFT and the permanent link to the model and data may be found at (?, ?). The raw ionosonde data may be found at https://giro.uml.edu/didbase/. The raw SSN, Kp, and F10.7 indices were obtained from the GFZ German Research Center for Geosciences, https://kp.gfz.de/en/data, the direct link to the text file containing the historical record is found at https://kp.gfz.de/app/files/Kp_ap_Ap_SN_F107_since_1932.txt. Finally, raw Dst data were obtained through the WDC for Geomagnetism, Kyoto http://wdc.kugi.kyoto-u.ac.jp/wdc/Sec3.html. All of these data are also available in the fully processed dataset included in the permanent link (?, ?).
Conflict of Interest Statement
The authors have no conflicts of interest to disclose.
@@ Acknowledgments
The authors would like to thank Dr. Victoriya Forsythe for her assistance and input with running the PyIRI model, the code for which can be found at (?, ?). The authors would like to also thank Dr. Ivan Galkin from the LGDC for the data they continue to collect and make available. We would also like to acknowledge the authors of (?, ?, ?) for the code they made available and which initially formed a basis for constructing our model.